OpenStudy (anonymous):

PLEASE HELP!!!!!!!!!!!!! If s = u + 0.5at^2, how would it be when solved for t?

OpenStudy (kagıtucak):

Leave the t alone s-u=0.5at^2 2(s-u)=at^2 2(s-u)/a)=t^2 t=root [2(s-u)/a]

OpenStudy (anonymous):

which one of these is the answer?

OpenStudy (kagıtucak):

The first one

OpenStudy (anonymous):

THANK YOU :)

OpenStudy (kagıtucak):

You're welcome :)